OpenStudy (anonymous):

Simplify the expression 7n - 4 + 3n I have no idea how to work these problems

OpenStudy (anonymous):

7n - 4 + 3n = (7+3)n - 4 = 10n -4

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

7n and 3n are like terms, (both have n), which means you can add them together just like 7 + 3 = 10 so 7n - 4 + 3n =10n -4 the 4 doesn't have an n so it remains.

OpenStudy (anonymous):

The -4 I meant

OpenStudy (anonymous):

Ok that helps a lot. Thanks

OpenStudy (anonymous):

I'm glad you got it your welcome :)

OpenStudy (anonymous):

So I could add 5k = 4k the same way?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

If there is an addition sign in between, yes, the same way.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks